# Antenna Coding Empowered by Pixel Antennas

### v1

@@ V-C MIMO Pixel Antenna System Performance
V-D Correlation of Antenna Coders in Codebook
It is worthwhile further investigating the codebook design for antenna coding given the enhanced channel gain and capacity of SISO/MIMO pixel antenna system optimized with codebook. To that end, we consider the correlation coefficient between the radiation patterns coded by the antenna coders in the codebook, defined as
$\rho_{i,j}=\frac{\mathbf{e}^{H}\left(\mathbf{b}_{i}\right)\mathbf{e}\left(\mathbf{b}_{j}\right)}{\left\|\mathbf{e}\left(\mathbf{b}_{i}\right)\right\|\left\|\mathbf{e}\left(\mathbf{b}_{j}\right)\right\|},$ (56)
where $\mathbf{b}_{i},\mathbf{b}_{j}\in\mathcal{B}$ are the antenna coders $i$ and $j$ in the codebook $\mathcal{B}$ and $\rho_{i,j}$ is the correlation coefficient between the radiation patterns coded by $\mathbf{b}_{i}$ and $\mathbf{b}_{j}$. The correlation coefficient, satisfying $\left|\rho_{i,j}\right|\in\left[0,1\right]$, can be used to evaluate how close the two patterns are, that is the two patterns are orthogonal when $\left|\rho_{i,j}\right|=0$ and are identical when $\left|\rho_{i,j}\right|=1$.
Moreover, the correlation coefficient $\rho_{i,j}$ can be used to characterize the channel correlation for the SISO pixel antenna system coded by different antenna coders. Specifically, according to (8), the channel for the SISO pixel antenna system, with the pixel antenna at the receiver coded by $\mathbf{b}_{i}$, is given by
$h\left(\mathbf{b}_{i}\right)=\mathbf{e}^{H}\left(\mathbf{b}_{i}\right)\mathbf{H}_{\textrm{V}}\mathbf{e}_{\textrm{T}},$ (57)
where $\left\|\mathbf{e}_{\textrm{T}}\right\|=1$, $\left\|\mathbf{e}\left(\mathbf{b}_{i}\right)\right\|=1$, and $\left[\mathbf{H}_{\textrm{V}}\right]_{i,j}$ $\forall i,j$ follows i.i.d. $\mathcal{CN}\left(0,1\right)$. Therefore, it is straightforward to show that given antenna coders $\mathbf{b}_{i}$ and $\mathbf{b}_{j}$ in the codebook, the channel correlation between $h\left(\mathbf{b}_{i}\right)$ and $h\left(\mathbf{b}_{j}\right)$ can be written as
$\mathbb{E}\left[h\left(\mathbf{b}_{i}\right)h^{\ast}\left(\mathbf{b}_{j}\right)\right]=\mathbb{E}\left[\mathbf{e}^{H}\left(\mathbf{b}_{i}\right)\mathbf{H}_{\textrm{V}}\mathbf{e}_{\textrm{T}}\mathbf{e}_{\textrm{T}}^{H}\mathbf{H}_{\textrm{V}}^{H}\mathbf{e}\left(\mathbf{b}_{j}\right)\right]=\rho_{i,j}.$ (58)
That is to say, when the SISO pixel antenna system switches the antenna coder from $\mathbf{b}_{i}$ to $\mathbf{b}_{j}$, the channel is switched from $h\left(\mathbf{b}_{i}\right)$ to $h\left(\mathbf{b}_{j}\right)$ with correlation $\rho_{i,j}$ given by (56).
The magnitude of the correlation coefficient between the radiation patterns coded by antenna coder 1 and the other antenna coders for different codebook sizes and physical apertures is shown in Fig. 11. From Fig. 11, we can find that the magnitude of the correlation coefficient is approximately uniformly distributed in $\left[0,1\right]$, especially when the codebook size is large, which means that the radiation patterns coded by other antenna coders in the codebook gradually deviate from the radiation pattern coded by antenna coder 1 until the radiation pattern becomes orthogonal. Actually, this is consistent with that the codebook is designed to provide as much diverse radiation patterns as possible to maximize the average channel gain. More importantly, as per (58), it is implied that when gradually switching antenna coders from 1 to others in the codebook, the channel for SISO pixel antenna system is accordingly switched with channel correlation gradually decreasing until the channel becomes independent. This is similar to the mechanism of switching “positions” in a fluid antenna system if we map the antenna coders in the codebook to “positions” in [17]. Nevertheless, it should be noted that the pixel antenna system with codebook design is more general and flexible with antenna configuration optimization fully considering the physical and EM property.
Figure 11: Magnitude of correlation coefficient between the radiation patterns coded by antenna coder 1 and the other antenna coders for different codebook sizes with physical apertures (a) $0.25\lambda\times 0.25\lambda$ and (b) $0.5\lambda\times 0.5\lambda$.
VI Conclusions

### v2

@@ V-C MIMO Pixel Antenna System Performance
V-D Evaluation of Multiport Circuit Network Model
To validate the accuracy of the multiport circuit network model, we compare it with the full-wave EM simulation using CST studio suite. Specifically, in CST studio suite, we simulate the pixel antenna with physical aperture $0.5\lambda\times 0.5\lambda$ configured by different antenna coders in the codebook (with codebook size of 4). The radiation patterns of the pixel antenna, obtained by the EM simulation and the multiport circuit network model (6) and (7), are shown in Fig. 11. We can find that the EM simulated radiation patterns are almost same as those computed by the multiport circuit network model for different antenna coders, showing the accuracy of the proposed model. In addition, the radiation patterns for different antenna coders are diverse, covering different angle ranges, which is consistent with the fact that the codebook is designed to maximize the ergodic channel gain in a multipath propagation environment with 2-D uniform power angular spectrum.
More importantly, under the same accuracy, the multiport circuit network model is significantly more computationally efficient. We measure the computational time to calculate the radiation pattern using the EM simulation (CST studio suite) and the proposed multiport circuit network model, respectively, in a laptop (with AMD Ryzen 9 7940HX CPU and 16 GB memory). The average computational time for the EM simulation and the multiport circuit network model, denoted as $T_{\textrm{EM}}$ and $T_{\textrm{circuit}}$, is measured around 56.59s and 0.000535s, respectively, which shows that the proposed model is around $10^{5}$ time faster than the EM simulation.
[…]
Last but not least, we completely compare the computational time for the whole process to optimize the antenna coder using the proposed multiport circuit network model and the EM simulation in the following two cases.
First, when using SEBO to optimize the antenna coder, i.e. solving the problem (25)-(26), the total number of calculations of radiation pattern is $N_{\textrm{e}}2^{J}$, so the total computational time for the proposed multiport circuit network model is
$T_{\textrm{total,circuit}}=T_{0}+N_{\textrm{e}}2^{J}T_{\textrm{circuit}},$ (69)
while the total computational time using EM simulation is
$T_{\textrm{total,EM}}=N_{\textrm{e}}2^{J}T_{\textrm{EM}}.$ (70)
If considering $N_{\textrm{e}}=4$ (the least number to optimize all the $Q$ switches), we have $T_{\textrm{total,circuit}}=\textrm{722}\,\textrm{s}$ and $T_{\textrm{total,EM}}=\textrm{231792}\,\textrm{s}.$
Second, when using the codebook to optimize the antenna coder, we need to design the codebook, which needs $N_{\textrm{c}}=N_{\textrm{o}}\left(\sum_{m=1}^{M}N_{\textrm{e},m}2^{J}\right)$ calculations of radiation pattern, where $N_{\textrm{o}}$ is the number of iterations of antenna coder optimization in GLA and $N_{\textrm{e},m}$ is the number of iterations for using SEBO to optimize the $m$th antenna coder in codebook with size of $M$. Once the codebook is ready, we search the codebook by $M$ calculations of radiation pattern. Totally, we need $N_{\textrm{c}}+M$ calculations of radiation pattern to use codebook to optimize antenna coder, so the total computational time for the proposed multiport circuit network model is
$T_{\textrm{total,circuit}}=T_{0}+\left(N_{\textrm{o}}\left(\sum_{m=1}^{M}N_{\textrm{e},m}2^{J}\right)+M\right)T_{\textrm{circuit}},$ (71)
while the total computational time using EM simulation is
$T_{\textrm{total,EM}}=\left(N_{\textrm{o}}\left(\sum_{m=1}^{M}N_{\textrm{e},m}2^{J}\right)+M\right)T_{\textrm{EM}}.$ (72)
If considering $N_{\textrm{o}}=1$ (the least number to perform antenna coding optimization in GLA), $N_{\textrm{e},m}=4$ $\forall m$, and e.g. $M=16$, we have $T_{\textrm{total,circuit}}=\textrm{755}\,\textrm{s}$ and $T_{\textrm{total,EM}}=\textrm{3709587}\,\textrm{s}$.
From the above complete comparison, we have shown that the proposed multiport circuit network model is significantly more computationally efficient than the EM simulation.
Figure 11: Radiation patterns of the pixel antenna (with physical aperture $0.5\lambda\times 0.5\lambda$) configured by antenna coders in the codebook (with codebook size of 4) using EM simulation and multiport circuit network model. Unit: dB.
VI Discussions and Future Work
VI-A Continuous Antenna Coding
It is possible to continuously adjust the load impedance $z_{\textrm{L},q}$ to implement a continuous antenna coding. To that end, we can use variable reactive load, such as varactor, to replace the RF switch in the pixel antenna, so that $z_{\textrm{L},q}$ can be modeled as $z_{\textrm{L},q}=jx_{\textrm{L},q}$, where $j$ is the imaginary unit and $x_{\textrm{L},q}$ is the continuous variable reactance. Accordingly, the continuous antenna coding design becomes a continuous optimization problem, which is easier to solve and can provide better performance compared with the binary antenna coding. However, implementing the continuous antenna coding increases the circuit complexity to control the pixel antenna since continuous dc bias voltages are required. Therefore, there is a tradeoff between the circuit complexity, system performance, and difficulty for optimization.
VI-B Antenna Coding Optimization for More Scenarios
Antenna coding can be optimized for more scenarios in future research, including, but are not limited to, the following:
VI-B1 Multi-User MIMO Pixel Antenna System
We can jointly optimize the precoding, combining, and antenna coding at transmitter and users to enhance the signal-to-interference-plus-noise ratio to maximize the weighted sum rate.
VI-B2 OFDM MIMO Pixel Antenna System
We can derive the frequency dependent model for pixel antenna and subsequently jointly optimize the precoding at each carrier frequency and the antenna coding of pixel antenna to maximize the capacity.
VI-B3 Pixel Antenna System for Wireless Power Transfer
We can jointly optimize the waveform, beamforming, and antenna coding to maximize the output dc power.
VI-C Pixel Antenna Element Spacing
There is a limitation on the spacing between adjacent pixel antenna elements, which should be neither too small nor too large, as explained in the following.
VI-C1 Inter-Element Electromagnetic Interference
When the spacing is too small, mutual coupling between the pixel antennas will be high and degrade the performance. In this work, we address the inter-element electromagnetic interference (EMI) issue in the MIMO pixel antenna system through separating the pixel antennas by a suitable distance to suppress the mutual coupling. For example, we separate two pixel antennas with physical aperture $0.25\lambda\times 0.25\lambda$ and $0.5\lambda\times 0.5\lambda$ by half-wavelength and one wavelength, respectively. Moreover, we can extend the proposed multiport circuit network model to take the inter-element EMI issue into account. For example, we consider two closely placed pixel antennas, which can be modeled as a ($2Q+2$)-port circuit network characterized by a $(2Q+2)\times(2Q+2)$ impedance matrix $\mathbf{Z}$. Given the antenna coders for pixel antennas 1 and 2, denoted as $\mathbf{b}_{1}$ and $\mathbf{b}_{2}$, we can find the corresponding load impedance $\mathbf{Z}_{\textrm{L},1}\left(\mathbf{b}_{1}\right)$ and $\mathbf{Z}_{\textrm{L},2}\left(\mathbf{b}_{2}\right)$. Thus, from $\mathbf{Z}$, $\mathbf{Z}_{\textrm{L},1}\left(\mathbf{b}_{1}\right)$, and $\mathbf{Z}_{\textrm{L},2}\left(\mathbf{b}_{2}\right)$, we can find the currents $\mathbf{i}_{1}\left(\mathbf{b}_{1},\mathbf{b}_{2}\right)$ and $\mathbf{i}_{2}\left(\mathbf{b}_{1},\mathbf{b}_{2}\right)$ and radiation patterns $\mathbf{e}_{1}\left(\mathbf{b}_{1},\mathbf{b}_{2}\right)$ and $\mathbf{e}_{2}\left(\mathbf{b}_{1},\mathbf{b}_{2}\right)$ for pixel antennas 1 and 2. Using the beamspace channel model, we can link the radiation patterns to the channel, so that we can optimize $\mathbf{b}_{1}$ and $\mathbf{b}_{2}$ to enhance the performance. By this way, the inter-element EMI issue can be fully characterized and considered. However, the extended model implies that the radiation pattern of individual pixel antenna is jointly coded by all the antenna coders due to the EMI issue, making the model complicate and intractable and increasing the complexity for antenna coding optimization.
VI-C2 Spherical Wave
When the spacing is too large, it makes the MIMO pixel antenna system very large, so the assumption of plane wave that we use becomes inaccurate. In this work, we assume that pixel antenna element spacing is not too large, as mentioned above, so that the plane wave assumption is valid. Otherwise, the effect of spherical wave should be incorporated in beamspace channel model for MIMO pixel antenna systems.
VI-D Pixel Antenna System Prototyping
Various pixel antennas have been designed, prototyped, and experimented. For example, in [8] and [10], the experiments for the pixel antenna prototypes have verified that the pixel antennas can flexibly reconfigure the radiation pattern toward different directions. Considering a line-of-sight (LOS) channel scenario, such pixel antennas can be directly used to generate an adaptive beam toward the receiver to enhance the system performance such as channel gain. Thus, the experiments provided in [8] and [10] have preliminarily verified the advantages of pixel antenna system in a simple LOS scenario. However, the previous work [8] and [10] were investigated at the level of antenna hardware design and limited to LOS scenario. To overcome this limitation, in this work we propose antenna coding which considers the impact on wireless communication with multipath fading at the system level to enhance the performance. Thus, in future research, the proposed pixel antenna system can be prototyped similarly to [8] and [10], but needs more efforts in the signal processing to control the pixel antenna with the proposed antenna coding technique.
VII Conclusions